How long will Php 60,000 accumulate to Php 62, 430 if the simple interest rate agreed is 1.50%?


1
Expert's answer
2021-11-24T15:12:59-0500

"A=P(1+rt)"


"62420=60000(1+0.015t)"


"t=(62420\/60000-1)\/0.015=2.69" years